OpenStudy (thomas5267):

Find all pairs of points with integer coordinates in the x-y plane which lie on the line 3x+5y=13.

ganeshie8 (ganeshie8):

By inspection, can you find ANY one solution ?

OpenStudy (thomas5267):

One particular solution is \(x=1,\,y=2\). \[ \begin{align*} 3(1+p)+5(2+q)&=13\\ 3+3p+5+5q&=13\\ 3p+5q&=0\\ p=-5k,&\,q=3k\\ x=1-5k,&\,y=2+3k \end{align*} \] I originally tried to solve this and it doesn't make sense. \[ 3(1+p)+5(1+q)=13 \]

ganeshie8 (ganeshie8):

You're almost done right after finding one particular solution : (x, y) = (1, 2)

ganeshie8 (ganeshie8):

x = 1-5k y = 2+3k is the general solution! which part doesn't make sense ?

OpenStudy (thomas5267):

I substitute the wrong number and it does not make sense lol. Now it does!

OpenStudy (thomas5267):

This makes me feel like that I am fishing for medals lol.

ganeshie8 (ganeshie8):

lol no, diophantine equations are not so easy to make sense of when you initially start...

ganeshie8 (ganeshie8):

If you notice , the same story repeats in number theory, linear algebra and differential equations : General solution = `particular solution` + `null solution` which is kind of amazing..

OpenStudy (thomas5267):

This is because we are studying with linear Diophantine equation, linear differential equation and linear algebra?

ganeshie8 (ganeshie8):

Maybe.. i dont knw exactly but i think this `particular` + `null` stuff shows up in many other areas that involve solving equations..

ganeshie8 (ganeshie8):

here is another almost same way to look at your earlier solution : \[3x+5y=13\] By inspection \((x, y) = (1, 2)\) is a particular solution to the given eqn. And the null solution is given by \(t(-5, 3)\) because \( 3(-5t) + 5(3t) = 0\) So Genral solution = `particular solution` + `null solution` = \((1, 2) + t(-5, 3)\)

ganeshie8 (ganeshie8):

You don't need to do any work for finding the null solution as above trick works in general. The null solution of equaiton \(ax+by = c\) is \(t(-b, ~a)\). So all that matters in solving a linear diophantine equation is finding the particular solution.

OpenStudy (thomas5267):

That's one heck of a insight.